# A Unified Bilevel Model for Adversarial Learning and A Case Study

### v1

@@ 3 Perturbation Analysis for Convex Clustering
In this section, we take clustering as an example to study the effect of perturbation. Specifically, due to the strongly convex nature of convex clustering model as well as the theoretical recovery result in [31], we choose the convex clustering model (CVC) in Example 2 as our learning model. We start with the case where the small perturbation will not change the clustering result.
@@
[31, Theorem 5] Consider the input data $X=\left[{x}_{1},\cdots,{x}_{n}\right]\in\mathbb{R}^{d\times n}$ and its partitioning $\mathcal{V}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$. Assume that all centroids $\left\{{x}^{(1)},{x}^{(2)},\ldots,{x}^{(K)}\right\}$ are distinct. Let $q\geq 1$ be the conjugate index of $p$ such that $\frac{1}{p}+\frac{1}{q}=1$. $Y^{*}$ is the unique solution of (CVC) and define the map $f\left({x}_{i}\right)={y}_{i}^{*}$ for $i=1,\cdots,n$. Let
@@
Theorem 1 shows that if conditions (C1) and (C2) hold, then the clustering result $D^{CVC}_{Y^{*}}$ coincides with the grand truth partition of $X$. That is, $D^{CVC}_{Y^{*}}=\mathcal{V}$. One can see that the exact recovery is based on conditions (C1) and (C2). Moreover, (C1) and (C2) are calculated based on the ground truth partition $\mathcal{V}$ as well as the data $X$. Given the perturbed data $X(\varepsilon)$, one can make use of Theorem 1 and provide a sufficient condition under which the clustering result is unchanged under perturbation $\varepsilon$. To that end, let $X(\varepsilon)=[x_{1}(\varepsilon),\cdots,x_{n}(\varepsilon)]$ be the perturbation of $X\in\mathbb{R}^{d\times n}$ with $\varepsilon\in\mathbb{R}^{d\times n}$. Under the partition of $D^{CVC}_{Y^{*}}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$, we define the following notations
@@
Here we use ${x}^{(\alpha)}({\varepsilon}),\ w^{(\alpha,\beta)}({\varepsilon})$ amd $w_{i}^{(\beta)}({\varepsilon})$ to mean that those coefficients may be related to the perturbation $\varepsilon$.
@@
$w_{ij}(\varepsilon)>0\ \text{and}\ n_{\alpha}w_{ij}(\varepsilon)>u_{ij}^{(\alpha)}(\varepsilon)\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,\dots,K.$ (C1)
@@
Proof. By applying [31, Theorem 5 ] with the ground truth partitioning $\mathcal{V}=D^{CVC}_{Y^{*}}$, and $X$ replaced by $X(\varepsilon)$, we get that the mapping $f_{\varepsilon}:X(\varepsilon)\rightarrow Y^{*}(\varepsilon)$ recovers the partitioning $\mathcal{V}$. That is, $D^{CVC}_{Y^{*}(\varepsilon)}$ is the same as $D^{CVC}_{Y^{*}}$. The proof is finished. $\square$
We demonstrate this by the following one-dimension example with two clusters, that is, $d=1$ and $K=2$. We uses the weighted matrix $W=(w_{ij})=E_{n}$ ( $E_{n_{1}}$ denotes the matrix of size $n_{1}\times n_{1}$ whose elements are all ones) and $p=2$. We only consider adding perturbation to a specific data. The solution of (CVC) is obtained by running the algorithm semismooth Newton-CG augmented Lagrangian method (Ssnal)  [^2] https://www.polyu.edu.hk/ama/profile/dfsun//Codes/Statistical-Optimization/ in [31].
@@
Figure 1: X=[0, 2, 10, 14]X=[0,\ 2,\ 10,\ 14]
We perturb only on $x_{3}$. Let $X(\varepsilon)=[0,\ 2,\ 17,\ 14]$, that is, $\varepsilon=[0,\ 0,\ 7,\ 0]$. Easily see that $\mu_{12}^{(1)}(\varepsilon)=0,\ \mu_{34}^{(2)}(\varepsilon)=0$. Moreover, $n_{\alpha}w_{ij}(\varepsilon)-u_{ij}^{(\alpha)}(\varepsilon)=n_{\alpha}>0,\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,2$, and $\gamma_{\min}^{\varepsilon}=\max\left\{\frac{2}{2},\frac{3}{2}\right\}=\frac{3}{2}<\gamma_{\max}^{\varepsilon}=\frac{14.5}{4}=\frac{29}{8}$. Therefore, conditions (C1) and (C2) hold and $D^{CVC}_{Y^{*}(\varepsilon)}=D^{CVC}_{Y^{*}}$. In fact, the solution of (CVC) gives $Y^{*}(\varepsilon)=[1,\ 1,\ 15.5,\ 15.5]$.
@@
Figure 2: $x_{3}(\varepsilon)\in\left(\frac{44}{5},\frac{68}{3}\right)$
@@
Figure 3: $X(\varepsilon)=[0,\ 2,\ -4,\ 14]$
@@ 4 Bilevel Model for Adversarial Learning
In this part, we will reformulate adversarial learning by bilevel optimization models. For general modeling and algorithmic background on bilevel optimization, see [42, 43].
As we mentioned in Section 3, the perturbation $\varepsilon$ on $X$ may led to the changed decision function, which means that the learning model is robust. On the other hand, the perturbation $\varepsilon$ on $X$ may lead to the change in decision result $D_{Y^{*}(\varepsilon)}$. It has two implications. Firstly, it means that such kind of noise is not neglectable. One needs to either do the denoising process to get rid of such noise or to improve the training process to make the learning model more robust. Secondly, from data attacking point of view, attack happens in such situation. In this case, a question arises: how could we choose the perturbation $\varepsilon$ such that some attacking criteria is maximized or minimized? This leads to the following two models for adversarial learning. The first model is to make the output of the perturbed learning model as huge difference as possible compared to the original learning model, which is described as ($a>0$ is given)
$\begin{split}\max_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon),\\ &\ \|\varepsilon\|\leq a,\end{split}$ (BL1)
@@
    $U\left(\varepsilon\right)=0$. That is, if there is no perturbation ($\varepsilon$=0), the deviation should be zero.
In (BL1), $\|\varepsilon\|\leq a$ controls the magnitude of $\varepsilon$ such that it would not be too large, otherwise, the perturbation will be recognized, and the attack will fail. One can see that (BL1) is a bilevel optimization problem where $Y^{*}(\varepsilon)\in S(\varepsilon)$ describes the lower level problem, saying that $Y^{*}(\varepsilon)$ must be the solution of (Pε).
@@
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ \|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon),\\ &\ U\left(\varepsilon\right)\geq\delta_{0}.\end{split}$ (BL2)
Based on the two base bilevel models (BL1) and (BL2), one can consider some variants for adversarial learning. One typical example is to combine the deviation function and the scale of $\varepsilon$ by a penalty function, that is,
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)+\rho\|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon).\end{split}$
By choosing different weight $\rho>0$, one can balance the deviation function and the scale of the perturbation according to the user’s preference.
For example, the bilevel model for convex clustering is given as follows
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)+\rho\|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)=\arg\min_{Y\in\mathcal{Y}}L^{CVC}(X(\varepsilon),Y)\end{split}$
Since $L^{CVC}(\cdot,\cdot)$ is given, so it is white-box adversaial attack. How to solve this bilevel problem heavily depends on the choices of the deviation function $U(\cdot)$.
Here we would like to highlight the following two important issues.
  • (i)
    Firstly, due to different learning models, $U(\cdot)$ is usually difficult to design. As far as we know, there is little work focusing on developing efficient deviation function, which is in fact important to the two adversarial learning models. We will address this question in the scenario of clustering, which can be found in Section 5.
  • (ii)
    Secondly, due to the potential discontinuity of decision function $D$ and the possibly complicated learning models, it is usually difficult to solve the adversarial models (BL1) or (BL2), which is out of the scope of this paper and will be further discussed in our future work.
@@ 5 A Case Study on Deviation Functions
In this part, we will study the $\delta$-measure function in asversarial learning for clustering problems, to verify whether it is a deviation function.
@@
For clustering problems, recall the aim of clustering is to partition the data points into different groups, i.e., $D^{CVC}_{Y^{*}(\varepsilon)}$ is a partition of points in $X(\varepsilon)$, there are many different ways of measuring the clustering results. See [44, 45] for some of the clustering functions. One natural way is to use a matrix to represent the partition of points. Take $\mathcal{V}=\{V_{1},\cdots,V_{K}\}$ as an example. A 0-1 matrix $\widehat{D}(X)\in\mathbb{R}^{n\times K}$ is defined as follows:
@@
The following function is proposed by Biggio et al. [46]
@@
Chhabra et al. [5] believed that $\delta$ increases with the number of points that spill over from partition $V_{1}$ to $V_{2}$ for $K=2$. However, it is still not quite clear about whether this function can fully represent the deviation of $D_{Y^{*}(\varepsilon)}$ over the original $D_{Y^{*}}$. Therefore, below we conduct a systematic analysis on the property of $\delta$ defined in (2). We consider the following two scenarios.

### v2

@@ 3 Perturbation Analysis for Convex Clustering
In this section, we take clustering as an example to study the effect of perturbation. We choose the convex clustering model due to the following reasons. Firstly, the strong convexity of $L^{CVC}(\cdot)$ guarantees the unique solution of (CVC), that is, $\mathcal{S}(\varepsilon)$ is a singleton for each $\varepsilon$. Secondly, there are exact recovery theoretical results under proper assumptions, which states that under some conditions, the solution of (CVC) perfectly recovers the ground truth clustering. Finally, Ssnal [sun2021convex] was proposed to solve (CVC) which is proved to be highly efficient. We will use it to solve (CVC) in our numerical experiments. We start with the case where the small perturbation will not change the clustering result.
@@
[sun2021convex, Theorem 5] Consider the input data $X=\left[{x}_{1},\cdots,{x}_{n}\right]\in\mathbb{R}^{d\times n}$ and its partitioning $\mathcal{V}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$. Assume that all centroids $\left\{{x}^{(1)},{x}^{(2)},\ldots,{x}^{(K)}\right\}$ are distinct. Let $q\geq 1$ be the conjugate index of $p$ such that $\frac{1}{p}+\frac{1}{q}=1$. $Y^{*}$ is the unique solution of (CVC) and define the map $f\left({x}_{i}\right)={y}_{i}^{*}$ for $i=1,\cdots,n$. Let
@@
Theorem 1 shows that if conditions (C1) and (C2) hold, then the clustering result $D^{CVC}_{Y^{*}}$ coincides with the ground truth partition of $X$. That is, $D^{CVC}_{Y^{*}}=\mathcal{V}$. One can see that the exact recovery is based on conditions (C1) and (C2). Moreover, (C1) and (C2) are calculated based on the ground truth partition $\mathcal{V}$ as well as the data $X$. Given the perturbed data $X(\varepsilon)$, one can make use of Theorem 1 and provide a sufficient condition under which the clustering result is unchanged under perturbation $\varepsilon$. To that end, let $X(\varepsilon)=[x_{1}(\varepsilon),\cdots,x_{n}(\varepsilon)]$ be the perturbation of $X\in\mathbb{R}^{d\times n}$ with $\varepsilon\in\mathbb{R}^{d\times n}$. Under the partition of $D^{CVC}_{Y^{*}}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$, we define the following notations
@@
Here we use ${x}^{(\alpha)}({\varepsilon}),\ w^{(\alpha,\beta)}({\varepsilon})$ and $w_{i}^{(\beta)}({\varepsilon})$ to mean that those coefficients may be related to the perturbation $\varepsilon$.
@@
$w_{ij}(\varepsilon)>0\ \text{and}\ n_{\alpha}w_{ij}(\varepsilon)>\mu_{ij}^{(\alpha)}(\varepsilon)\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,\dots,K.$ (C1)
@@
Proof. By applying [sun2021convex, Theorem 5 ] with the ground truth partitioning $\mathcal{V}=D^{CVC}_{Y^{*}}$, and $X$ replaced by $X(\varepsilon)$, we get that the mapping $f_{\varepsilon}:X(\varepsilon)\rightarrow Y^{*}(\varepsilon)$ recovers the partitioning $\mathcal{V}$. That is, $D^{CVC}_{Y^{*}(\varepsilon)}$ is the same as $D^{CVC}_{Y^{*}}$. The proof is finished. $\square$
We demonstrate this by the following one-dimension example with two clusters, that is, $d=1$ and $K=2$. We use the weighted matrix $W=(w_{ij})=E_{n}$ ( $E_{n_{1}}$ denotes the matrix of size $n_{1}\times n_{1}$ whose elements are all ones) and $p=2$. We only consider adding perturbation to a specific data. The solution of (CVC) is obtained by running the algorithm semismooth Newton-CG augmented Lagrangian method (Ssnal)  [^2] https://www.polyu.edu.hk/ama/profile/dfsun//Codes/Statistical-Optimization/ in [sun2021convex].
@@
Figure 1: The original one-dimensional dataset X=[0, 2, 10, 14]X=[0,\ 2,\ 10,\ 14] to illustrate convex clustering and its robustness to data perturbations.
We perturb only on $x_{3}$. Let $X(\varepsilon)=[0,\ 2,\ 17,\ 14]$, that is, $\varepsilon=[0,\ 0,\ 7,\ 0]$. Easily see that $\mu_{12}^{(1)}(\varepsilon)=0,\ \mu_{34}^{(2)}(\varepsilon)=0$. Moreover, $n_{\alpha}w_{ij}(\varepsilon)-\mu_{ij}^{(\alpha)}(\varepsilon)=n_{\alpha}>0,\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,2$, and $\gamma_{\min}^{\varepsilon}=\max\left\{\frac{2}{2},\frac{3}{2}\right\}=\frac{3}{2}<\gamma_{\max}^{\varepsilon}=\frac{14.5}{4}=\frac{29}{8}$. Therefore, conditions (C1) and (C2) hold and $D^{CVC}_{Y^{*}(\varepsilon)}=D^{CVC}_{Y^{*}}$. In fact, the solution of (CVC) gives $Y^{*}(\varepsilon)=[1,\ 1,\ 15.5,\ 15.5]$.
@@
Figure 2: Perturbed dataset $X(\varepsilon)$ with $x_{3}(\varepsilon)\in\left(\tfrac{44}{5},\tfrac{68}{3}\right)$, where the conditions (C1) and (C2) hold and the clustering results remain unchanged.
@@
Figure 3: Perturbed dataset $X(\varepsilon)=[0,\ 2,\ -4,\ 14]$, where condition (C2) fails and the convex clustering result changes from $\left\{\{1,2\},\{3,4\}\right\}$ to $\left\{\{1,2,3\},\{4\}\right\}$.
@@
Here we would like to highlight the following two important issues.
  • (i)
    The assumptions of distinct centroids and strictly positive weights, which appear in Theorem 2, are inherited from the recovery result of convex clustering [sun2021convex, Theorem 5] and are standard in that context. These conditions ensure that the convex clustering model admits a unique, well-separated partition and allow us to derive clean sufficient conditions under which this partition is preserved under perturbations. For practical clustering problems, these assumptions may fail, putting the exact recovery in question. Therefore, it would be interesting to further study the exact recovery result when the above assumptions fail. We leave this interesting question as our future research topic to study.
  • (ii)
    As mentioned in Section 2.2, the perturbation of $X(\varepsilon)$ can be in various forms. The results in Theorem 2 and Corollary 1 hold not only for the additive point-wise noise, but also hold for other noise situations, such as multiplicative noise, feature-level noise.
@@ 4 Bilevel Model for Adversarial Learning
In this part, we will reformulate adversarial learning by bilevel optimization models. For general modeling and algorithmic background on bilevel optimization, see [kleinert2021survey, chen2025set].
As we mentioned in Section 3, the perturbation $\varepsilon$ on $X$ may lead to the changed decision function, which means that the learning model is robust. On the other hand, the perturbation $\varepsilon$ on $X$ may lead to the change in decision result $D_{Y^{*}(\varepsilon)}$. It has two implications. Firstly, it means that such kind of noise is not neglectable. One needs to either do the denoising process to get rid of such noise or to improve the training process to make the learning model more robust. Secondly, from data attacking point of view, attack happens in such situation. In this case, a question arises: how could we choose the perturbation $\varepsilon$ such that some attacking criteria is maximized or minimized? This leads to the following two models for adversarial learning. The first model is to make the output of the perturbed learning model as huge difference as possible compared to the original learning model, which is described as ($a>0$ is given)
$\begin{split}\max_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in\mathcal{S}(\varepsilon),\\ &\ \|\varepsilon\|\leq a,\end{split}$ (BL1)
@@
    $U\left(0\right)=0$. That is, if there is no perturbation ($\varepsilon=0$), the deviation should be zero.
In (BL1), the constraint $\|\varepsilon\|\leq a$ controls the magnitude of $\varepsilon$, so that the perturbation is not too large. Otherwise, it would be easily detected and the attack would fail. In practice, the threshold $a$ can be calibrated using prior knowledge on admissible noise levels in a given application. For example, many clustering robustness studies explicitly test algorithms under injected noise ratios around $15\%$ to $30\%$, such as adding $30\%$ uniform or Gaussian noise to the data [li2007noise, zhang2016large], and one may choose $a$ in accordance with such empirical noise levels. One can see that (BL1) is a bilevel optimization problem where $Y^{*}(\varepsilon)\in\mathcal{S}(\varepsilon)$ describes the lower level problem, saying that $Y^{*}(\varepsilon)$ must be the solution of (Pε).
@@
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ \|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in\mathcal{S}(\varepsilon),\\ &\ U\left(\varepsilon\right)\geq\delta_{0}.\end{split}$ (BL2)
Remark 2.
$U(\cdot)$ is usually difficult to design due to different learning models. As far as we know, there is little work focusing on developing efficient deviation function, which is in fact important to the two adversarial learning models. We will address this question in the scenario of clustering, which can be found in Section 5.
Bilevel models (BL1) and (BL2) are flexible to accommodate many concrete learning tasks. At the lower level, one may take $\ell_{2}$-regularized least squares, logistic regression, support vector machines, or deep neural networks [franceschi2018bilevel, mackay2019hyperparameter]. The deviation function $U(\cdot)$ may take prediction loss on a validation set or robust loss in adversarial training [madry2018towards, zhang2022fastat]. Moreover, $U(\cdot)$ may also represent robustness-oriented criteria inspired by risk-driven anti-clustering, such as the risk-weighted diversity objective used to partition training data in [mauri2023riskanticlustering]. Exploring these deviation choices within our bilevel models thus offers several promising directions for future research.
Various approaches can be explored to solve the adversarial model (BL1). For the white-box setting, where the lower-level model is fully accessible, many approaches can be employed, such as the KKT-based approaches [dempe2002foundations, colson2007overview], value-function-based approaches [dempe2013value, gao2022value], the duality-based approaches [ouattara2018duality, li2023novel] and the hypergradients-based approaches [andrychowicz2016learning, maclaurin2015gradient, finn2017model]. On the other hand, the bilevel model also applies to black-box attacks. In such cases, the lower-level model does not admit explicit mathematical formula, one may use evolutionary strategies [wierstra2014natural, hansen2016cma] or derivative-free methods [zhang2024prima, li2011class] to solve (BL1).
Comparing (BL2) with (BL1), (BL2) is usually more challenging to solve. The reason is that the feasible region of $U\left(\varepsilon\right)\geq\delta_{0}$ is difficult to explicitly represent due to the implicit form of $Y^{*}(\varepsilon)$. Therefore, how to design efficient algorithms to solve (BL2) is an interesting topic which is worth further investigation.
For convex clustering problem, bilevel models (BL1) and (BL2) reduce to the following form
[…]
We will discuss how to choose $U(\cdot)$ for clustering in the following section.
@@ 5 A Case Study on Deviation Functions
In this part, we will study the $\delta$-measure function in adversarial learning for clustering problems, to verify whether it is a deviation function.
@@
For clustering problems, recall the aim of clustering is to partition the data points into different groups, i.e., $D^{CVC}_{Y^{*}(\varepsilon)}$ is a partition of points in $X(\varepsilon)$, there are many different ways of measuring the clustering results. See [yin2024rapid, feng2023review] for some of the clustering functions. One natural way is to use a matrix to represent the partition of points. Take $\mathcal{V}=\{V_{1},\cdots,V_{K}\}$ as an example. A 0-1 matrix $\widehat{D}(X)\in\mathbb{R}^{n\times K}$ is defined as follows:
@@
The following function is proposed by Biggio et al. [biggio2013data]
@@
Chhabra et al. [chhabra2020suspicion] believed that $\delta$ increases with the number of points that spill over from partition $V_{1}$ to $V_{2}$ for $K=2$. However, it is still not quite clear about whether this function can fully represent the deviation of $D_{Y^{*}(\varepsilon)}$ over the original $D_{Y^{*}}$. Therefore, below we conduct a systematic analysis on the property of $\delta$ defined in (3). We consider the following two scenarios.